OpenStudy (anonymous):

NeED HELPP!!!!! Jack and Jill went shopping for school supplies. Jack bought 4 pencils and 2 pens for $2.60. Jill bought 2 pencils and 5 pens for $3.30. Write a systems of eqautions to represent their orders. a. How much does a pencil cost? b. How much does a pen cost?

OpenStudy (anonymous):

\[4\times pencils +2\times pens =2.6\\2\times pencils +5\times pens =3.3\]

OpenStudy (anonymous):

can I use X and Y variables to represent pens and pencils versus writing the word

OpenStudy (anonymous):

??

OpenStudy (anonymous):

That should be fine, but just make sure to clearly state which is which. For example x = pens y = pencils or something like that :)

OpenStudy (anonymous):

cool thnx

OpenStudy (anonymous):

No prob